A particle executes a simple unrestricted random walk on the line: a step to the right of length 1 occurs with probability 0.56 and a step to the let of length 1 occurs with probability 0.44. Initially the particle is at the onigin. Select the option that gives an approximate value for the probability that the particle will be less than 6 units from the origin after 100 steps, calculated using an appropriate continuity correction. Select one: 0.195 0.238 0.280 Ο 0.720 0.762 0.805
Answer #1

let x=1 with p=0.56 and x=-1 with p=0.44

here from above given data

x f(x) xP(x) x2P(x)
1 0.56 0.560 0.560
-1 0.44 -0.440 0.440
total 0.120 1.000
E(x) =μ= ΣxP(x) = 0.1200
E(x2) = Σx2P(x) = 1.0000
Var(x)=σ2 = E(x2)-(E(x))2= 0.9856
std deviation=         σ= √σ2 = 0.9928

therefore for 100 steps ; expected distance Y=100*E(X)=100*0.12=12

and std deviation =sqrt(100*Var(X))=sqrt(100*0.9856)=9.93

hence  probability that the particle will be less than 6 units from the origin after 100 steps

=P(-5.5<X<5.5)=P((-5.5-12)/9.93<Z<(5.5-12)/9.93)=P(-1.76<Z<-0.65)=0.2578-0.0392=0.2186

0.2186 is approximately equal to 0.238

so option b is correct